OpenStudy (anonymous):

A=1 _h(b+c) for b i really need help with this one 2

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (shadowfiend):

Hmm. Are you trying to solve for b? And do you have values for A/h/c or do you just want it in terms of A, h, and c?

OpenStudy (anonymous):

ok here it is...i got b=2A-c

OpenStudy (shadowfiend):

gfa -- you seem to have missed `h'.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

I am just having a hard time with this class

OpenStudy (shadowfiend):

So the steps are: \[\begin{align} A &= \frac{1}{2}h(b + c)\\ 2A &= h(b+c)\\ \frac{2A}{h} &= b + c\\ \frac{2A}{h} - c &= b \end{align}\] First you multiply both sides by 2 to cancel out the 1/2. Then you divide both sides by h to cancel out the h on the right. Then you subtract c from both sides. Finally, b is alone :)

OpenStudy (anonymous):

okay so i did it right then i wasn't sure

OpenStudy (anonymous):

i have just having a horrible time with this Algebra class i had a 92 in now i have an 80

OpenStudy (shadowfiend):

You did? Awesome! :) Always feels good to hear someone else confirm it. Remember to click `Good Answer' next to answers that are helpful to you here on OpenStudy :)

OpenStudy (anonymous):

you are right shdow i missed it lol

OpenStudy (shadowfiend):

Feel free to ask any other questions. Hopefully that grade can go back up :) Algebra can be tough, but if you keep pushing it'll make more and more sense.

OpenStudy (anonymous):

I appreciate the help i have a test that i am studying for

OpenStudy (anonymous):

I need it to go back up i have a 3.5 gpa i can't loose that

OpenStudy (shadowfiend):

All right then, good luck on that test!

OpenStudy (anonymous):

sure you don't mind

OpenStudy (anonymous):

i just have couple hard ones that i an having a hard time with

OpenStudy (shadowfiend):

Yeah, I'd post them separately on the left, that way people can try to help with each one when they think they can.

OpenStudy (anonymous):

okay thanks i will give me a few minutes i dropped my calculator